OpenStudy (anonymous):

At which value will the graph of y=cotx have a zero? pi/2 pi/3 pi/4 none

OpenStudy (anonymous):

can you think of where tangent is undefined?

OpenStudy (anonymous):

0?

OpenStudy (whpalmer4):

\[\cot x = \frac{\cos x}{\sin x}\]Where does \(\cos x=0\)?

OpenStudy (anonymous):

no idea @whpalmer4

OpenStudy (whpalmer4):

Get out your unit circle diagram and look at it...

OpenStudy (whpalmer4):

Make yourself some flashcards and start learning those values.

OpenStudy (anonymous):

is pi/2 the answer @whpalmer4

OpenStudy (whpalmer4):

If \(\cos \frac{\pi}{2} = 0\) then it is the correct answer.

OpenStudy (whpalmer4):

(yes)